OpenStudy (anonymous):

A baseball is thrown across a horizontal surface and follows a path described by the equation x^(2)+25/16y^(2) -400x=0 (all dimensions are in feet). What is the highest point the ball reaches and how far from the starting point (in horizontal distance) does this happen? (What is the apex of the trajectory?)

OpenStudy (harsimran_hs4):

want to do it by quadratic eqs or calculus?

OpenStudy (anonymous):

calculus

OpenStudy (harsimran_hs4):

ok let's begin we need to find maximum value of y so differentiate the eq wrt to x and tell me

OpenStudy (anonymous):

I dont know. Im all lost with all this. thanks for your effort

OpenStudy (anonymous):

Im doing parabolas and hyperbolas equations but with this one idk what to do

OpenStudy (harsimran_hs4):

differentiating wrt to x \[2x + \frac{ 25 }{ 16 }(2y)\frac{ dy }{ dx } - 400 = 0\] now from this eq rearrange and get dy/dx on one side and other terms on different side can you?

OpenStudy (harsimran_hs4):

well let me tell you this is neither a eq of parabola or hyperbola so check it again

OpenStudy (anonymous):

\[2x-400=- \frac{ 25 }{ 16 } y^2\]

OpenStudy (harsimran_hs4):

do the same thing but we are taking horizontal distance so we need to maximize x so differentiate the eq wrt to x can you?

OpenStudy (anonymous):

\[2x=400 -\frac{ 25 }{ ? }\]

OpenStudy (anonymous):

\[2x=400 -\frac{ 25 }{ 16 } y^2\]

OpenStudy (anonymous):

\[x=\frac{ 6400-25y^2 }{ 8} ?????\]

OpenStudy (harsimran_hs4):

ok but i told to differentiate so differentiating we get \[2\frac{ dx }{ dy } = 2\frac{ 25 }{ 16 } y \]

OpenStudy (anonymous):

i dont get it. sorry

OpenStudy (harsimran_hs4):

fo what value of y is dx/dy maximum when y is 0 (figure out why)

OpenStudy (anonymous):

0

OpenStudy (harsimran_hs4):

so maximum value of x is 200